# Feature-Based Interpretable Optimization

### v1

@@ 5.3.2 Varying Grid Size and Features
Due to the type of graphs used, the reasonable choice of parameters $F$ and $n$ is limited. The results for experiments where those were varied are presented in Table 1. It can be seen that when increasing the size of the graph whilst using the same value for $F$, the performance of all methods except MICROx for which it drops, increases. The use of larger graphs furthermore results in larger computation times for all methods used. MIPx can be solved remarkably fast. For both extreme cases when varying $F$ described in detail in Section 5.2, the known and expected behavior can be recognized. In general, the performance on the training as well as on the test data decreases for all methods except MICROx. Except for the latter, the computation time needed increases with the number of used features. It is worth mentioning that the MIPx methods can be solved significantly faster for $F=81$ than MICROx, although both represent the same problem.
@@ 5.3.3 Chicago graph
For the trees generated using the Chicago graph, we restricted ourselves to trees with a maximum depth of two and the use of the instance (meta) features weekday and time. In Table 2 the scaled objectives and computational times are presented. Our example tree presented in Figure 4 is referred to as ART4.
@@
Training 0.99 0.87 0.90 0.90 0.99
Test 0.69 0.33 0.84 0.84 0.87
Time (s) 54.24 4.99 7.31 4.99
@@
It can be seen that congruent with the previous results, all the methods using meta-solutions achieve better objective values than our benchmark MICRO4. Especially, both of the heuristics lead to good results on the test data. Note that although ART4 archives the same objective as MIP4 this does not imply that this tree is optimal since we allowed more complex queries. Even though we were using a large graph, the methods MIP4 and MICRO4 could be solved in an acceptable time.
5.4 Discussion and Insights
@@
Summarizing the presented results, we see that solving MIPx results in genuinely good solutions. It was shown that for most instances the formulations could not be solved to optimality within the time limit of 900 seconds. Using this method therefore takes much more time than solving the nominal problems. However, it should be noted, that in real-world applications the found optimization rule would be used many times, which puts the computational effort into perspective. Furthermore, feasible solutions were found in every case within the time limit.
@@
Using the learning heuristic (LHx), we generated solutions that were often nearly as good, and in some cases even better, than those produced by MIPx. This is remarkable as the trees could be generated using the heuristic in a fraction of the time limit. Regarding the objective value, heuristic M2Mx did perform mostly worse than MIPx and LHx. Furthermore, it has been shown to not scale well due to the necessity of solving an MIP formulation as part of the process. This process can be sped up by using heuristics for generating trees with micro solutions assigned to their leaves, e.g. via the greedy heuristic presented in [GH23]. Our results imply LHx to be clearly preferable to M2Mx in general and to MIPx at least for big instances, which should occur in reality in most cases.
@@
Due to the restriction of only considering $K$ solutions and the requirements of the tree structure, none of our methods were able to achieve the objective of the black-box model (OPT). Nevertheless, in comparison to the interpretable approach from [GH23] (MICROx), our framework has the potential to produce – depending on the value of $F$ – better solutions. In some cases, trees with two leaves and meta-solutions even outperformed trees with four leaves and micro-solutions on the test data. Therefore we can further close the gap between the performance of interpretable methods and black-box models whilst being able to increase interpretability at the same time by the use of smaller trees.
@@
It could also be shown that an increasing ratio between the dimension of features and scenario size leverages the performance of our approach. For example, in the experiments presented in Section 5.2.3, for values of $F$ in the interval $[2,6]$ whilst keeping the instance size fix, MIP4 results in solutions which are 47.57% better on average compared to MICRO4 considering their scaled objective. For values of $F$ in $[7,15]$ a genuinely smaller gap of 25.57% on average – but in favor of MIP4 – can be observed. Without the context of a real-world application, our results suggest therefore to set $F$ to one or at least a relatively small value. A small value of $F$, though, increases the requirements on the user of the decision tree since the solution space described by the solution features increases. This could lead to a point where, in reality, optimal micro-solutions can not be guaranteed and the quality of solutions decreases. The ability to work with a given meta-solution is highly dependent on the user and the given use case. We therefore were not able to consider this in our computational experiments. The performance gain in real-world applications is hence related to a reasonable choice of the solution features and their dimensions.
@@
Our presented framework is suitable if an interpretable solution process for an optimization problem is needed and the person who is executing the generated optimization rules has some domain knowledge as well as the authorization to make decisions on their own. If the latter is not the case, we recommend using the approach of [GH23], where strict solutions are found during the optimization process. Both approaches have in common, that they require historical or at least estimated data for the generation of the optimization rules. Since both use small decision trees to represent/illustrate the optimization rules, interpretability is ensured which helps to increase user acceptance.
@@
Our approach provides optimization rules that map a given scenario to a space of possible solutions. It therefore allows – and requires – a final decision, which is the actual solution to implement. This more dynamic approach allows the user to react to unpredictable events and hence generate better solutions. Furthermore, it helps to avoid undesired micromanagement.
@@
The options for generating training data are case-specific. If meaningful historical data is available, its use can be considered, otherwise data must be estimated. Depending on the size of the problem, MIPx or LHx should be used for the training of the decision tree. For most bigger instances, applying LHx will be beneficial due to a much shorter runtime, whilst for small instances or in use cases where sufficient time is available, MIPx can potentially generate better solutions. The solution features have to be chosen in a way, which enables the user to utilize the meta-solutions and still provides a sufficient degree of freedom.
@@ 6 Conclusion
Interpretable optimization aims at bridging the gap between optimization theory and practice by taking the use of methods by decision-makers and stakeholders into account. By providing solution methods that are transparent and easy to apply, we avoid what is de facto black-box optimization in practice and therefore improve acceptance of optimization algorithms. As this topic has been introduced only recently, the current toolbox of interpretable algorithms is still very limited. In a previous work, a decision tree was used to map instance parameters to solutions.
@@
In this paper, we introduce a more flexible framework, which is based on mapping instance features to solution features instead. This means that a post-optimization step becomes necessary to convert the resulting solution features (also referred to as meta-solutions) to a single, feasible solution. In the context of a shortest path problem, for example, this may mean that a given sequence of districts must be converted to a path. The granularity of features should thus be modeled in a way that a decision-maker can do this step.
@@
We demonstrated how to apply this feature in different settings, where we explain the context of knapsack problems and shortest path problems in particular detail. To apply our framework, we introduced mixed-integer programming formulations as well as heuristics, in particular a learning heuristic that leverages existing fast and reliable methods to construct classification trees. We analyzed the complexity of our framework and showed that simple and natural special cases already turned out to be hard. In extensive computational experiments involving randomly generated data as well as real-world data, we analyzed the performance of our framework and found that the additional flexibility provided by the use of features can greatly improve the quality of the produced optimization rule, thus reducing the cost of interpretability, i.e., the difference to what is achievable by classic, non-interpretable solution methods.
@@
As interpretable optimization is a recent field of research, we believe many interesting further possibilities to study arise. One aspect is to analyze the impact if the decision maker is not able to choose an optimal solution that corresponds to a given meta-solution. We might define a probability distribution over the set of all solutions that correspond to a meta-solution to obtain an expected performance metric (i.e., a stochastic approach), or even consider a worst-case solution from this set (i.e., a robust approach). Uncertainty may not only be present in the conversion of meta-solutions to solutions but also in the problem data. In this case, it may be beneficial to construct decision trees that take uncertainty in training data into account.

### v3

@@ 5.3.2 Varying Grid Size and Features
Due to the type of graphs used, the reasonable choice of parameters $F_{S}$ and $n$ is limited. The results for experiments where those were varied are presented in Table 2. It can be seen that when increasing the size of the graph whilst using the same value for $F_{S}$, the performance of all methods—except MICROx—increases. The use of larger graphs furthermore results in larger computation times for all methods used. MIPx can be solved remarkably fast. At the extreme points, when varying $F_{S}$, as described in detail in Section 5.2, the known and expected behavior can be recognized. In general, the performance on the training as well as on the test data decreases for all methods except MICROx. Except for the latter, the computation time needed increases with the number of features used. It is worth mentioning that the MIPx methods can be solved significantly faster for $F_{S}=81$ than MICROx, although both represent the same problem.
@@ 5.3.3 Chicago Graph
For the trees generated using the Chicago graph, we restricted ourselves to trees with a maximum depth of two and the use of the instance (meta) features weekday and time. In Table 3, the scaled objectives and computational times are presented. Our example tree presented in Figure 5 is referred to as ART4.
@@
Training 0.95 0.90 0.92 0.94 0.92
Test 0.86 0.40 0.82 0.78 0.89
Time (s) 23.22 6.22 6.08 6.22
@@
It can be seen that, congruent with the previous results, all the methods using meta-solutions achieve better objective values than our benchmark MICRO4. Especially, both of the heuristics lead to good results on the test data. Even though we were using a large graph, the methods MIP4 and MICRO4 could be solved in an acceptable time.
5.4 Discussion and Managerial Insights
@@
Summarizing the presented results, we see that solving MIPx results in genuinely good solutions. It was shown that for most instances, the formulations could not be solved to optimality within the time limit of 900 seconds. Using this method, therefore, takes much more time than solving the nominal problems. However, it should be noted that in real-world applications, the found optimization rule would be used many times, which amortizes the computational effort. Furthermore, feasible solutions were found in every case within the time limit.
@@
Using the learning heuristic (LHx), we generated solutions that were often nearly as good, and in some cases even better, than those produced by MIPx. This is remarkable as the trees could be generated using the heuristic in a fraction of the time limit. Regarding the objective value, the heuristic M2Mx performed mostly worse than MIPx and LHx. Furthermore, it has been shown not to scale well due to the necessity of solving an MIP formulation as part of the process. This process can be sped up by using heuristics for generating trees with micro solutions assigned to their leaves, e.g., via greedy heuristics. Our results imply LHx to be preferable to M2Mx in general and to MIPx at least for big instances, which should occur in reality in most cases.
@@
Due to the restriction of only considering $K$ solutions and the requirements of the tree structure, none of our methods were able to achieve the objective of the black-box model (OPT). Nevertheless, in comparison to the interpretable approach from Goerigk and Hartisch (2023) (MICROx), our framework has the potential to produce—depending on the value of $F_{S}$—better solutions. In some cases, trees with two leaves and meta-solutions even outperformed trees with four leaves and micro-solutions on the test data. Therefore, we can further close the gap between the performance of interpretable surrogates and black-box models whilst being able to increase interpretability at the same time by the use of smaller trees.
@@
It could also be shown that an increasing ratio between the dimension of features and scenario size leverages the performance of our approach. For example, in the experiments presented in the Electronic Companion, for values of $F_{S}$ in the interval $[2,6]$ whilst keeping the instance size fixed, MIP4 results in solutions which are 47.57% better on average compared to MICRO4 considering their scaled objective. For values of $F_{S}$ in $[7,15]$ a smaller gap of 25.57% on average—but in favor of MIP4—can be observed. Without the context of a real-world application, our results suggest therefore to set $F_{S}$ to a relatively small value while ensuring that the selected features are still interpretable. A small value of $F_{S}$, though, increases the requirements on the user of the decision tree since the solution space described by the solution features increases. This could lead to a point where, in reality, optimal micro-solutions can not be guaranteed and the quality of solutions decreases. Note that this result does not hold in general. For example, there are cases in which adding meaningless features does not result in a decrease in solution quality and vice versa.
The ability to work with a given meta-solution is highly dependent on the user and the given use case. We therefore do not consider this in our computational experiments. The performance gain in real-world applications is hence related to a reasonable choice of the solution features and their dimensions.
@@
Our presented framework is suitable if an interpretable solution process for an optimization problem is needed, and the decision maker who is executing the generated optimization rules has some domain knowledge as well as the authorization to make decisions on their own. If the latter is not the case, we recommend using an interpretable method, where strict solutions are found during the optimization process. Both approaches require historical or at least estimated data for the generation of the optimization rules. Since both use small decision trees to represent the optimization rules, interpretability is ensured which helps to increase user acceptance.
@@
Our approach provides optimization rules that map a given scenario to a space of possible solutions. It therefore allows—and requires—a final decision, which is the actual solution to implement. This more dynamic approach allows the user to react to unpredictable events and hence generate better solutions. Furthermore, it helps to avoid undesired micromanagement.
@@
The options for generating training data are case-specific. If meaningful historical data is available, its use can be considered, otherwise, data must be estimated. Depending on the size of the problem, MIPx or LHx should be used for the training of the decision tree. For most bigger instances, applying LHx will be beneficial due to a much shorter runtime, whilst for small instances or in use cases where sufficient time is available, MIPx can potentially generate better solutions. The solution features have to be chosen in a way that enables the user to understand the meta-solutions while still providing a sufficient degree of freedom.
@@ 6 Conclusion
Interpretable optimization surrogates aim to bridge the gap between optimization theory and practice by accounting for the methods used by decision-makers and stakeholders. By providing solution methods that are transparent and easy to apply, we avoid what is de facto black-box optimization and therefore improve the acceptance of optimization algorithms. As this topic has been introduced only recently, the current toolbox of interpretable methods is still very limited.
@@
In this paper, we introduce a flexible framework to find the best optimization rule to map instance features to solution features. This means that a post-optimization step becomes necessary to convert the resulting solution features (also referred to as a meta-solution) to a single, feasible solution. In the context of a shortest path problem, for example, this may mean that a given sequence of districts must be converted to a path. The granularity of features should thus be modeled in a way that a decision-maker can perform this step.
@@
We demonstrated how to apply this framework in different settings, with a special focus on knapsack and shortest path problems. We introduced mixed-integer programming formulations as well as heuristics, in particular a learning heuristic that leverages existing fast and reliable methods to construct classification trees. We analyzed the complexity of our framework and showed that simple and natural special cases already turned out to be hard. In extensive computational experiments involving randomly generated data as well as real-world data, we analyzed the performance of our framework and found that the additional flexibility provided by the use of features can greatly improve the quality of the produced optimization rule, thus reducing the cost of interpretability, i.e., the difference to what is achievable by classic, non-interpretable solution methods. This reduced cost, along with the inherent benefits brought on by interpretability, such as acceptability and responsiveness, makes it worthwhile as a manager to put resources into exploring the benefits of interpretability and to investigate how different stakeholders might react to more comprehensible optimization results.
@@
As interpretability in optimization is a recent field of research, we believe there are many further directions to study. One aspect is to analyze the impact if the decision maker is not able to choose an optimal solution for a given meta-solution and instead relies on suboptimal outcomes. In this case, we might define a probability distribution over the set of all solutions that correspond to a meta-solution to obtain an expected performance metric (i.e., a stochastic approach), or even consider a worst-case solution from this set (i.e., a robust approach). Uncertainty may not only be present in the conversion of meta-solutions to solutions but also in the problem data. In this case, it may be beneficial to construct decision trees that take uncertainty in training data into account.
7 Code and Data Disclosure
The code and data to support the numerical experiments in this paper is available on demand.
Appendix A The Necessity of Preventing Circles in the Solution of Shortest Path Based Instances
The question of whether it is necessary to actively avoid cycles in the concept presented in Chapter 3.3 arises, since it seems counterintuitive that an optimal solution to a shortest path problem with non-negative edge costs contains cycles when optimizing w.r.t. the Laplace criterion. As the assumption of having to suppress them plays a significant role in the proofs presented in Appendix B, this question has direct and clear effects on the correctness of the hardness results. We therefore present an example illustrating the necessity of prohibiting them.
Figure 10: Example for a shortest path instance with two scenarios and two solution features highlighting the need to actively avoid cycles in the solutions.
Consider the instance of a shortest path problem depicted in Figure 10. We are given two cost scenarios which are displayed on the corresponding edges. Let $M$ be a large number. Furthermore, let $f(s)=f(t)=red$ and $f(1)=f(2)=blue$. We now want to determine a meta solution for finding a path from $s$ to $t$ such that the costs with respect to the Laplace criterion are minimized.
First, cycles in the paths are not considered prohibited. In this case, the (unique) optimal meta path is the sequence (red-blue-red) which allows the paths (s-1-s-t) for the first and (s-1-2-t) for the second scenario. This meta solution, therefore, results in an objective value of $6$. It becomes obvious that this solution exploits the possibility of cycles in the paths. If this meta solution is evaluated in an environment where paths containing cycles are considered infeasible, the optimal path in scenario one becomes (s-1-2-t), resulting in an overall objective value of $M+5$.
Indeed, the (also unique) optimal solution for the problem of finding an optimal meta path if cycles are prohibited is given by the sequence (redred) and the corresponding paths (s-t) and (s-t). This results in an objective value of only $M+1$.
Note that this example highlights the necessity of actively suppressing cycles in paths when solving the problem of finding the best meta solution, as well as when solving the problem of finding optimal paths for a given meta solution.
Appendix B Hardness Results
We now turn to the complexity of finding interpretable surrogates for the shortest path problem. As it is possible to solve the nominal shortest path problem in graphs without negative cycles efficiently, the question arises whether polynomial-time algorithms also exist for finding interpretable optimization rules in this case. In the following, we give negative answers to two special cases.
Theorem 1.
Proof.
Proof A reduction from the NP-complete Hamiltonian path problem (Garey and Johnson, 1979) is constructed. Given an undirected graph $G=(V,E)$ with nodes $s,t\in V$, we need to decide whether there is a (simple) $s$-$t$-path visiting all nodes in $V$ exactly once. We first create a graph $\bar{G}=(\bar{V},\bar{E})$ in the following way: For every $v\in V$, two nodes $v$ and $v^{\prime}$ are added to $\bar{G}$ as well as an edge $(v,v^{\prime})$. Node $v$ is assigned to district $a$, and node $v^{\prime}$ to district $b$. For every $\{i,j\}\in E$, the edges $(i^{\prime},j)$ and $(j^{\prime},i)$ are added to $\bar{G}$. Figure 11 shows an example of this construction.
Figure 11: Example of construction for the proof of Theorem 1.
Theorem 2.
The following problem is NP-complete: Given a graph $G=(V,E)$ where each node $v\in V$ belongs to a district $f(v)$, and a list of cost scenarios $\bm{c}^{1},\ldots,\bm{c}^{N}\in\{0,1\}^{E}$. Decide whether there is a meta path such that there exist paths that satisfy the meta path which have a cost of 0 for each scenario. In particular, the problem of finding a single meta-path with minimum costs is not approximable, even in directed acyclic graphs.
Proof.
Proof Let an instance of the NP-complete 3SAT problem (Garey and Johnson, 1979) be given. It consists of $n$ Boolean variables $x_{1},\ldots,x_{n}$ and a list of clauses $C_{1},\ldots,C_{m}$, where each clause is a disjunction of three literals. We need to decide if there is a truth assignment of the $n$ variables such that all $m$ clauses are true.
We first show how to construct a graph $G^{\prime}$ with $2n+2$ nodes contained in the set $V^{\prime}$. For each variable $x_{i}$, we construct a node $t_{i}$ and a node $f_{i}$, which represent the states of “true” and “false”, respectively. We further introduce nodes $s$ and $t$ to be used as the source and target nodes of the meta-path. The set of edges is given by
$\cup\left\{(s,t_{1}),(s,f_{1}),(t_{n},t),(f_{n},t)\right\}.$
$E=$ $E_{1}\cup E_{2}$
$\cup\{(t^{1}_{i},t^{2}_{i}):i\in[n]\}\cup\{(f^{1}_{i},f^{2}_{i}):i\in[n]\}$
$\cup\{(s,t^{1}_{1}),(s,f^{1}_{1}),(t^{2}_{n},t),f^{2}_{n},t)\}$
$\text{with }E_{1}=$ $\Big\{(a,b):a\in\{t^{1}_{i},f^{1}_{i}\},\ b\in\{t^{1}_{i+1},f^{1}_{i+1}\},\ i=1,\ldots,n-1\Big\}$
Finally, we introduce $m$ cost vectors $\bm{c}^{j}\in\{0,1\}^{E}$, each of them corresponding to a clause $C_{j}$. The costs for all edges in $E_{1}$ and $E_{2}$ are always zero. The cost of an edge $(t^{1}_{i},t^{2}_{i})$ is zero if $x_{i}$ is contained in $C_{j}$, otherwise, the costs are one. Analogously, the cost of an edge $(f^{1}_{i},f^{2}_{i})$ is zero if $\bar{x}_{i}$ is contained in $C_{k}$, otherwise, the costs are one. Finally, the costs of edges $(s,t^{1}_{1})$, $(s,f^{1}_{1})$, $(t^{2}_{n},t)$, and $(f^{2}_{n},t)$ are always zero.
Figure 13 illustrates the construction of $G$ with the cost scenario that corresponds to the clause $(x_{1}\vee\bar{x}_{2}\vee x_{3})$. There are only three edges with cost one, indicated by a dotted arrow. These are the edges $(f^{1}_{1},f^{2}_{1})$, $(t^{1}_{2},t^{2}_{2})$, and $(f^{1}_{3},f^{2}_{3})$.
Figure 13: Graph in proof of Theorem 2. Edges represented by solid lines have cost of zero, these represented by dotted lines have cost one.
Appendix C Content to Be Included in the Electronic Companion
In this document, the results of extensive experiments on instances of knapsack problems are presented. All experiments were conducted with respect to the setting described in Sections 5.1 and 5.2.
C.0.1 Varying Number of Items
First, we present experiments in which the number of items was varied. For these, the number of training scenarios $N$ was set to ten, and the number of solution features $F_{S}$ was set to four. Figures 14(a) and 14(b) illustrate the scaled objective value on the training data versus the number of items provided per instance. In Figure 14(a), the solid green and red lines are partially obscured by the blue line. It can be observed that there is a negative correlation between the number of items and the objective for all trees with two and four leaves. This trend is more distinct for trees with two leaves. Moreover, the methods that generate trees with four leaves demonstrate considerably better performance. However, it is remarkable that those from LH4 perform significantly worse than the other methods, including the benchmark. This can be explained by the fact that the trees are constructed with regard to the accuracy and not the objective function value of the knapsack problems. As illustrated in Figure 14(b), the performance of LH2 and LH4 as well as the performance of M2M2 increases again as the number of items rises. Even with the performance of M2M2 increasing, for 140 or more items, this method is not able to outperform META1. This means that for these instances, the tree structure cannot be used beneficially. Moreover, the quality of the solutions from MICRO2 and MICRO4 continues to decline. Notably, the performance differences between MICRO2 and M2M2 are considerably larger than those between MICRO4 and M2M4.
(a) objective training/ #items.
(b) objective training/ #items.
(c) objective test/ #items.
(d) objective test/ #items.
(e) runtime/ #items.
(f) runtime/ #items.
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption $M2$
Refer to caption META1
Figure 14: Plots for knapsack and synthetic data, #items on x-axis. A higher scaled objective indicates a better result. The solution-based approach, which serves as a benchmark, is highlighted in red.
Figures 14(c) and 14(d) depict the relationship between the scaled test objective and the number of items per instance. For all methods except MICRO2 and MICRO4, a positive correlation can be identified. For instances with a small number of items, the methods MIP4, MICRO4, M2M4 are overfitting. The gap between MICROx and M2Mx derived from it increases significantly with an increasing number of items. In contrast to its performance on the training data, the performance on the test data of LH2 and LH4 is highly competitive. For instances with many items, significantly better performance is achieved when using META1 than MICROx. In general, it can be concluded that the performance of our methods appears to be clearly advantageous when evaluated on test data.
The median computing time used is shown in Figures 14(e) and 14(f). The time required for solving via M2Mx results almost exclusively from the time used to calculate MICROx. The data points of MICROx, therefore, also represent the computing time of M2Mx. Even small instances for MIP4 could not be solved to optimality without hitting the time limit of 900 seconds. It can even be seen that all MIP-based methods except MICRO2 cannot be solved to optimality within the time limit. In every run, feasible solutions were found within the time limit. LHx can be solved for all configurations within seconds. For both MICRO2 and MICRO4, dips in the progressions can be observed.
C.0.2 Varying Number of Scenarios
Figures 15(e) and 15(f) show the computation time for a varying number of scenarios with $n=16$ and $F_{S}=4$. For MIP4, even instances with ten scenarios couldn’t be solved to optimality. Also, when using MIP2 and MICRO4, the time limit was hit at 16 and 18 training scenarios respectively. The computation times for MICRO2 show a notable dip from 14 to 23 scenarios, which we were not able to explain. In contrast to the MIP-based approaches, when using LH2 and LH4, we can generate decision trees in less than two seconds even for instances with 200 scenarios.
(a) objective training/ #scenarios.
(b) objective training/ #scenarios.
(c) objective test/ #scenarios.
(d) objective test/ #scenarios.
(e) runtime/ #scenarios.
(f) runtime/ #scenarios.
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption $M2$
Refer to caption META1
Figure 15: Plots for knapsack and synthetic data, #scenarios on x-axis. A higher scaled objective indicates a better result. The solution-based approach, which serves as a benchmark, is highlighted in red.
The relationship between the scaled objective on the training set and the number of scenarios used for training is shown in Figures 15(a) and 15(b). For all methods generating trees with two leaves, the influence on the training performance is marginal. The performance of MICRO4 and M2M4 decreases significantly when more than 60 scenarios are used, and even falls below that of MICRO2 and M2M2. This can be explained because no (proven) optimal solutions could be determined for MICRO4. It can be seen that the quality of the solutions of MICRO4 and M2M4 is strongly correlated and that of M2M4 presumably depends on MICRO4.
In Figure 15(c), it can be seen that for all methods utilized, the performance increases with the number of scenarios used for training. In contrast, Figure 15(d) shows that the performance of MICRO4 and M2M4 begins to decline at a certain point as the number of scenarios increases. As with the data in Fig. 15(b), this is due to the intractability of the models. It is noteworthy that LH4 clearly outperforms MIP4. In general, all developed methods outperform the benchmark model of the same size. Furthermore, it can be seen that the marginal benefit of further training scenarios decreases.